Two people, one in Los Angeles and one in New York, take the graduate record entrance exam. These were different exams, but we would like to compare their scores. The person in Los Angeles received a 647 while the average 640 with a standard deviation of 20. The person in New York recieved a 570 while the avarage was 560 with a standard deviation of 25. First, calculate the z- score for each person, then indicate who scored higher after standing the scores. Person 2 (new york)??
Answer #1

z score for Person 1 = (647 - 640)/20 = 7/20 = 0.35

z score for Person 2 = (570 - 560)/25 = 0.40

Since z score for Person 2 i.e. person from New York is higher, Person 2 scored higher after standing the scores.